# Flavor-Changing Non-Global Logarithms

### v2

@@ II Calculating Non-Global Logarithms as a series in the Jet Radius
As mentioned above, because we restrict the soft quarks to not lie in the same jet, there are no collinear divergences. Additionally, there is no strongly-ordered soft limit because quark number is conserved. Therefore, the simplest possible observable that regulates the soft divergence is measuring the energy of one of the quark or anti-quark’s transverse momentum. As we have focused on a jet of interest $J$, we will ignore everything outside of it, and measure the transverse momentum of the soft particle that lies in the jet. The observable $\tau$ is then
@@ II.1 Correlations between the jet of interest and another jet
Here, $A$ denotes any other jet in the event. Note that the integration measure scales like $R^{4}$, so expansion through this order ensures terms through an overall scaling of $R^{2}$ are captured. The leading collinear matrix element is rather simple, where
@@ II.2 Correlations between two other jets
Note that now, there is no contribution at order-$R^{-4}$ because other jets in the event are assumed to lie an order-1 angle from one another. Additionally, while this expansion is simple enough to take, the resulting matrix element is extremely unwieldy, and I have not been able to find a compact expression for it. As such, I will not write it down here. Nevertheless, its integral over phase space does simplify nicely, where
@@
On the second line, the result has been expressed in a coordinate-independent way, while on the third line, I have not found a coordinate-independent way to express it, and so explicit $\eta,\phi$ dependence is presented. Here, $\Re$ denotes the real part.
@@
By contrast, I was unable to evaluate the integral over ${\cal I}_{AB}(\tilde{p}_{1}=n_{J},\tilde{p}_{2})$ with the native routines in Mathematica. However, studying the integral more closely, it can be evaluated. Let’s write down the relevant integral, where
@@
and perform the contour integral about the unit circle. The non-analytic structure inside the unit circle is rather simple in this case. First, there is only a single pole located at $z_{*}=\min[e^{\eta_{2}},e^{-\eta_{2}}]$, which comes from the overall factor $1/(n_{J}\cdot\tilde{p}_{2})^{2}$. Next, there are two branch cuts from the logarithms $\log(n_{A}\cdot\tilde{p}_{2})$ and $\log(n_{B}\cdot\tilde{p}_{2})$. These branch cuts live on the line segments $z_{\text{$A$,cut}}\in\left[0,e^{i\phi_{A}}\min[e^{\eta_{2}-\eta_{A}},e^{-\eta_{2}+\eta_{A}}]\right]$ and $z_{\text{$B$,cut}}\in\left[0,e^{i\phi_{B}}\min[e^{\eta_{2}-\eta_{B}},e^{-\eta_{2}+\eta_{B}}]\right]$, respectively. The non-analytic structure and contour on the $z$ plane we consider are displayed in Fig. 2. Evaluating the corresponding residues or integrals along the branch cuts is straightforward, and can be done analytically. The integral over $\eta_{2}$ can then be correspondingly evaluated; however, again, the final expression we find is horribly unwieldy and I have not been able to simplify it, so we do not report it here.
@@ III.1 A Subtractive Jet Flavor Algorithm
In step 2., note that individual partonic energy cuts are not collinear safe, but this is a bit of a distraction, and not a fundamental limitation. In particular, because we take the $p_{\perp,\text{cut}}\to 0$ limit, collinear divergences vanish at the end, but may arise in intermediate steps. To eliminate them, one can introduce an appropriate fragmentation function that absorbs potential collinear divergences. We construct the necessary fragmentation functions at one-loop order in App. A. However, we also note that the fragmentation functions trivialize for $p_{\perp,\text{cut}}\to 0$.
@@ III.2 Subtractive Jet Flavor at Fixed-Order
To fully validate and test the subtractive flavor algorithm requires complete predictions through next-to-next-to-leading order. While there are now several studies with these modern flavor algorithms at this order [49, 50, 51, 52, 53], we will restrict an analysis here to a sufficiently simple flavor configuration that illustrates the subtleties with non-global logarithms, but does so only at next-to-leading order. Specifically, we will consider the process $pp\to W^{+}+j$ at the 13 TeV LHC, and focus on the jet flavor $f_{j}=(bu)$, a composite flavor in which there is both a sufficiently hard bottom and down quark. This jet flavor configuration is chosen so that there is a single non-global logarithmic contribution, from the final state $W^{+}bu\bar{u}$. In particular, note that the final state $W^{+}b\bar{b}u$ is not possible at tree level which simplifies our analysis here.
@@ IV Looking Towards an All-Orders Subtractive Flavor Algorithm
$\sigma^{\text{(2)}}_{q\bar{q}\text{,NGL}}(p_{\perp,\text{cut}})$ $=-2\,C_{A}\,n_{f}\,T_{R}\,\sigma_{pp\to g}^{(0)}\left(\frac{\alpha_{s}}{2\pi}\right)^{2}\,\log\frac{\mu}{p_{\perp,\text{cut}}}\left(-\frac{2}{3}+\frac{4}{9}\pi^{2}\right)\,.$
@@
Our assumptions actually enable us to very simply resum the contribution from an arbitrary number of unresolved quarks or anti-quarks. In this limit, all soft quark–anti-quark pairs are distinct flavor, independent, and emitted with an identical distribution on phase space. Therefore, in the strongly-ordered limit and to all-orders, this cross section for gluon flavor with unresolved quarks split across the jet boundary simply exponentiates:
@@
where (so) denotes the strongly-ordered limit. Next, according to our prescription of subtractive jet flavor, we must subtract the non-global logarithm contribution to all orders. With these approximations, the non-global contribution is

### v3

@@ II Calculating Non-Global Logarithms as a series in the Jet Radius
As mentioned above, because we restrict the soft quarks to not lie in the same jet, there are no collinear divergences. Additionally, there is no strongly-ordered soft limit because quark number is conserved. Therefore, the simplest possible observable that regulates the soft divergence is to measure the energy of one of the quark or anti-quark’s transverse momentum. As we have focused on a jet of interest $J$, we will ignore everything outside of it, and measure the transverse momentum of the soft particle that lies in the jet. The observable $\tau$ is then
@@ II.1 Correlations between the jet of interest and another jet
Here, $A$ denotes any other jet in the event. Note that the integration measure scales like $R^{4}$, so expansion through this order ensures terms through an overall cross section scaling of $R^{2}$ are captured. The leading collinear matrix element is rather simple, where
@@ II.2 Correlations between two other jets
Note that now, there is no contribution at order-$R^{-4}$ because other jets in the event are assumed to lie an order-1 angle from one another. Additionally, while this expansion is simple enough to take, the resulting matrix element is extremely unwieldy, and we have not been able to find a compact expression for it. As such, we will not write it down here. Nevertheless, its integral over phase space does simplify nicely, where
@@
On the second line, the result has been expressed in a coordinate-independent way, while on the third line, we have not found a coordinate-independent way to express it, and so explicit $\eta,\phi$ dependence is presented. Here, $\Re$ denotes the real part.
@@
By contrast, we were unable to evaluate the integral over ${\cal I}_{AB}(\tilde{p}_{1}=n_{J},\tilde{p}_{2})$ with the native routines in Mathematica. However, studying the integral more closely, it can be evaluated. Let’s write down the relevant integral, where
@@
and perform the contour integral about the unit circle. The non-analytic structure inside the unit circle is rather simple in this case. First, there is only a single pole located at $z_{*}=\min[e^{\eta_{2}},e^{-\eta_{2}}]$, which comes from the overall factor $1/(n_{J}\cdot\tilde{p}_{2})^{2}$. Next, there are two branch cuts from the logarithms $\log(n_{A}\cdot\tilde{p}_{2})$ and $\log(n_{B}\cdot\tilde{p}_{2})$. These branch cuts live on the line segments $z_{\text{$A$,cut}}\in\left[0,e^{i\phi_{A}}\min[e^{\eta_{2}-\eta_{A}},e^{-\eta_{2}+\eta_{A}}]\right]$ and $z_{\text{$B$,cut}}\in\left[0,e^{i\phi_{B}}\min[e^{\eta_{2}-\eta_{B}},e^{-\eta_{2}+\eta_{B}}]\right]$, respectively. The non-analytic structure and contour on the $z$ plane we consider are displayed in Fig. 2. Evaluating the corresponding residues or integrals along the branch cuts is straightforward, and can be done analytically. The integral over $\eta_{2}$ can then be correspondingly evaluated; however, again, the final expression we find is horribly unwieldy and we have not been able to simplify it, so we do not report it here.
@@ III.1 A Subtractive Jet Flavor Algorithm
In step 2., note that individual partonic energy cuts are not collinear safe, but this is a bit of a distraction, and not a fundamental limitation. In particular, because we take the $p_{\perp,\text{cut}}\to 0$ limit, collinear divergences vanish at the end, but may arise in intermediate steps. To eliminate them, one can introduce an appropriate fragmentation function that absorbs potential collinear divergences. We construct the necessary fragmentation functions at one-loop order in App. A. However, we also note that the fragmentation functions trivialize for $p_{\perp,\text{cut}}\to 0$. Additionally, we note that such intermediate collinear divergences first arise at next-to-next-to-next-to-leading order (NNNLO) in perturbation theory, where there is a gluon emission, in addition to the soft quark–anti-quark pair. At this order, there is a single collinear divergence in this final state configuration when the gluon is collinear to the soft quark that lies in the jet. Therefore, the one-loop fragmentation functions can be used to subtract the collinear divergences that arise when imposing a finite $p_{\perp,\text{cut}}$. At higher orders, fragmentation functions at higher loops will be necessary to subtract collinear divergences.
We additionally note that subtractive jet flavor is very much a fixed-order flavor definition. Specifically, by definition, possible large logarithms are explicitly subtracted. This subtraction means that there are no large logarithms that may spoil perturbative convergence and require all-orders resummation. Therefore, potential collinear divergences in intermediate steps in the calculation of the subtractive jet flavor cross section can be dealt with on a case-by-case basis, order-by-order in perturbation theory that one calculates. As mentioned above, this is first an issue at NNNLO, which is beyond current state-of-the-art precision calculations relevant for jet physics.
In a practical implementation of this procedure, one would typically evaluate the jet flavor as a function of $p_{\perp,\text{cut}}$, for a range of small but finite and non-zero values, in a numerical fixed-order code. We present a first analysis of this procedure in Sec. III.2 using MadGraph, but it is worth commenting on numerical limitations here. At next-to-next-to-leading order, when this flavor definition is first non-trivial, there are only single soft logarithms that need to be subtracted to extract the finite limiting value as $p_{\perp,\text{cut}}\to 0$. This is therefore much simpler, and correspondingly much more numerically stable, than other cases in which fixed-order codes are used to extract finite residual contributions. For example, high-precision fixed-order codes have long been used to extract two-loop constants in factorization theorems for infrared and collinear safe observables, e.g., for thrust in Ref. [49]. However, in the case of the thrust observable for example, even the next-to-leading order cross section as the value of thrust goes to 0 is a fourth-order polynomial in the logarithm of the thrust, and so a substantially more delicate cancellation is necessary to extract the finite contribution. In Sec. III.2, we demonstrate that using the default settings in MadGraph can produce reasonable extractions of subtractive jet flavor cross sections, even with no attempt at numerical optimization.
@@ III.2 Subtractive Jet Flavor at Fixed-Order
To fully validate and test the subtractive flavor algorithm requires complete predictions through next-to-next-to-leading order. While there are now several studies with these modern flavor algorithms at this order [50, 51, 52, 53, 54], we will restrict an analysis here to a sufficiently simple flavor configuration that illustrates the subtleties with non-global logarithms, but does so only at next-to-leading order. Specifically, we will consider the process $pp\to W^{+}+j$ at the 13 TeV LHC, and focus on the jet flavor $f_{j}=(bu)$, a composite flavor in which there is both a sufficiently hard bottom and up quark. This jet flavor configuration is chosen so that there is a single non-global logarithmic contribution, from the final state $W^{+}bu\bar{u}$. In particular, note that the final state $W^{+}b\bar{b}u$ is not possible at tree level which simplifies our analysis here.
@@ IV Looking Towards an All-Orders Subtractive Flavor Algorithm
$\sigma^{\text{(2)}}_{q\bar{q}\text{,NGL}}(p_{\perp,\text{cut}})$ $=-2\,C_{A}\,n_{f}\,T_{R}\,\sigma_{pp\to g}^{(0)}\left(\frac{\alpha_{s}}{2\pi}\right)^{2}\,\log\frac{\mu}{p_{\perp,\text{cut}}}\left(-\frac{2}{3}+\frac{4}{9}\pi^{2}\right)\,.$ (78)
@@
Our assumptions actually enable us to very simply resum the contribution from an arbitrary number of unresolved quarks or anti-quarks. With these approximations and in the limit in which all soft quark–anti-quark pairs are emitted off of the hard jet strongly-ordered in energy and collinear to the jet direction, we can easily write down a differential equation that describes the evolution of the non-global quark flavor, as a function of $p_{\perp,\text{cut}}$. As $p_{\perp,\text{cut}}$ is decreased by a small amount $\delta p_{\perp}$, there are two possibilities: either the net quark flavor is unchanged or a new soft quark pair is emitted (or newly resolved) that is split across the boundary. As gluons carry no flavor the net quark flavor is completely blind to arbitrary gluon emission, and so resolved gluons do not appear in the evolution equation. Similarly, if both quarks in a soft $q\bar{q}$ pair stay in the jet or leave the jet, the net quark flavor is also unchanged. That is, for both gluon emission and soft quark pairs that both land in (out of) the jet, there is exact real-virtual cancellation and no large logarithms are generated.
These observations motivate first the finite difference equation for the strongly-ordered gluon jet flavor cross section $\sigma^{\text{(so)}}_{g}(p_{\perp,\text{cut}})$, where
$\sigma^{\text{(so)}}_{g}(p_{\perp,\text{cut}}-\delta p_{\perp})=\sigma^{\text{(so)}}_{g}(p_{\perp,\text{cut}})-\delta p_{\perp}\,\frac{2C_{A}n_{f}T_{R}\left(\frac{\alpha_{s}}{2\pi}\right)^{2}\left(-\frac{2}{3}+\frac{4}{9}\pi^{2}\right)}{p_{\perp,\text{cut}}}\,\sigma^{\text{(so)}}_{g}(p_{\perp,\text{cut}})\,.$ (79)
The first term on the right represents no new quark flavor resolved as $p_{\perp,\text{cut}}$ is decreased, while the second term describes the gluon flavor cross section lost by a newly resolved non-global $q\bar{q}$ pair in the interval $p_{\perp,\text{cut}}-\delta p_{\perp},p_{\perp,\text{cut}}]$. This follows from differentiating Eq. 78. The solution of the resulting differential equation for gluon flavor with unresolved quarks split across the jet boundary is
@@
where $\sigma_{pp\to g}^{(0)}$ is the cross section for gluon jet flavor at the high scale $\mu$. Next, according to our prescription of subtractive jet flavor, we must subtract the non-global logarithm contribution to all orders. With these approximations, the non-global contribution is
@@ V Conclusions
In a realistic experimental analysis, of course partons are not observable; only hadrons are. Further, hadron flavor is in general very challenging to identify at a modern collider, with bottom (or possibly charm) the exception. Like other recent jet flavor definitions, subtractive jet flavor would be most practically useful for precision studies of heavy flavor, specifically for jets in which bottom hadrons are present. Perturbatively, non-global soft quarks spoil the IRC safety of naïve jet flavor, but at the hadronic level, such soft quarks would be bound in color-singlet hadrons. Further, unlike perturbation theory with massless quarks, hadrons all have finite masses and must have non-zero momentum to be detected. Because of this, there are no non-global divergences of flavor at hadron level, and so there is nothing to subtract. [^5] Another recent jet flavor proposal is similar in this way to subtractive jet flavor on hadronized events [62]. This suggests that the implementation of subtractive jet flavor on hadronized events is identical to the implementation of naïve jet flavor, to simply sum together the net flavors of the hadrons in the jet, as all observed hadrons would satisfy the $p_{\perp,\text{cut}}\to 0$ cut. In some previous studies, naïve flavor jets have been shown to be very similar at the hadron level to fully IRC safe flavor definitions [28, 29]. More study of fixed-order, resummed, and hadronized predictions of subtractive jet flavor are necessary, but this observation may provide justification for why naïve jet flavor seems to work reasonably well in practice.